A group of 50 persons attended a secret meeting in a room 12m x 10m x 3m high. The

room is completely sealed off and insulated. Each person gives off 150 kcal/hr of heat

and occupies a volume of 0.2 cubic meters. The room has an initial pressure of 101.3

kpa and 16C. calculate the room temp after 10minutes.


1
Expert's answer
2022-03-03T13:23:00-0500

The volume of the air in the room is the difference between the volume of the room and the total volume occupied by people:


"V = 12\times 10\times 3 m^3 - 50\times 0.2m^3 = 350m^3"


Assuming the density of air to be "\rho = 1.225kg/m^3", the mass of the air is "m = \rho V".

The amount of heat given in 10 minutes is:


"Q = 150kcal/hr\times 1/6hrs\times 50 = 1250kcal"

The temperature difference is then:


"\Delta T = \dfrac{Q}{C_Vm} = \dfrac{Q}{c_v\rho V}"

where "c_V = 0.171 kcal\/kg\/K" is the specific heat capacity. Thus, obtain:


"\Delta T = \dfrac{1250kcal}{0.171kcal/kg/K\cdot 1.225kg/m^3\cdot 350m^3} \approx 17\degree C"

Thus, the final temperature:


"T_2 = 17\degree C + 16\degree C = 33\degree C"

Answer. "33\degree C".